# How does the optimizer implicitly bias the model merging loss landscape?

### v1

@@ Abstract
Model merging methods combine models with different capabilities into a single one while maintaining the same inference cost. Two popular approaches are linear interpolation, which linearly interpolates between model weights, and task arithmetic, which combines task vectors obtained by the difference between finetuned and base models. While useful in practice, what properties make merging effective are poorly understood. This paper explores how the optimization process affects the loss landscape geometry and its impact on merging success. We show that a single quantity – the effective noise scale – unifies the impact of optimizer and data choices on model merging. Across architectures and datasets, the effectiveness of merging success is a non-monotonic function of effective noise, with a distinct optimum. Decomposing this quantity, we find that larger learning rates, stronger weight decay, smaller batch sizes, and data augmentation all independently modulate the effective noise scale, exhibiting the same qualitative trend. Unlike prior work that connects optimizer noise to the flatness or generalization of individual minima, we show that it also affects the global loss landscape, predicting when independently trained solutions can be merged. Our findings broaden the understanding of how optimization shapes the loss landscape geometry and its downstream consequences for model merging, suggesting the possibility of further manipulating the training dynamics to improve merging effectiveness.
@@ 1 Introduction
Model merging combines independently trained solutions to fuse their unique capabilities into a single solution. These methods have been applied to either further improve the performance of a single model (Wortsman et al., 2022) or to combine models with different, but similar, capabilities (Ilharco et al., 2023). Notably, this is achieved without additional computational cost at inference time. Given these practical advantages, merging methods have been applied to improve state-of-the-art architectures (Yadav et al., 2024). However, these approaches rely on extensive trial-and-error, requiring practitioners to train and evaluate multiple models to identify which candidates successfully merge. A fundamental unresolved question is why some models are compatible for merging while others fail, despite achieving similar individual performance.
@@
The success of model merging depends on the mode connectivity phenomenon, the geometric property that independent solutions can be connected by paths of low-loss models (Draxler et al., 2018; Garipov et al., 2018)Frankle et al. (2020) demonstrated a stricter condition, linear mode connectivity, showing that solutions sharing a common initial optimization trajectory can be connected by a linear path with similar loss. This finding suggests that optimization dynamics, not just final convergence, fundamentally shape the geometry of the loss landscape between solutions. Yet, despite this insight, well-studied factors that affect the optimization dynamics, such as learning rate, weight decay, or batch size, are not well understood for the model merging loss landscape, thereby ignoring their potential role in determining merging compatibility.
@@
In this work, we study the role of optimization dynamics on the outcome of model merging. First, we show that different optimizer components (learning rate, weight decay, batch size, and data augmentation) control the same underlying factor, the effective noise scale, which captures the magnitude of stochastic fluctuations in the optimization trajectory. Our experiments demonstrate how the noise magnitude determines whether independently trained solutions occupy merging-compatible regions of the loss landscape. A noise regime that is too small or too large leads to almost no merging benefits, whereas a moderately large noise regime maximizes the benefits. After that, we decompose the noise into individual components, showing that each one exhibits the same qualitative trend. Starting from the learning rate component, we find that a larger (and stable) value can consistently identify solutions that merge more effectively than those trained with smaller learning rates, even when both converge to similar individual performance levels. Since decaying the learning rate (large or small) during the optimization can always lead to a converged solution, it is perhaps surprising that simply starting with a larger learning rate can change the merging outcomes. However, beyond classical research showing direct advantage of large learning rates on generalization (Keskar et al., 2016), recent works presented different implicit biases of training with a larger learning rate, such as a sparser activation (Andriushchenko et al., 2023b), a different sequence of pattern learning (Li et al., 2019), and a flatter solution (Andriushchenko et al., 2023a). Our results extend these benefits beyond single-task performance, showing that a larger learning rate unlocks more effective merging, quantified using performance gain (i.e. performance difference between the merged and the single model).
@@
Similarly to the learning rate, we find that a large weight decay also enables more effective merging, beyond improvements in the single model performance. We explain this phenomenon using the notion of the effective learning rate (Van Laarhoven, 2017; Hoffer et al., 2018), which suggests that the main role of weight decay is to prevent the (effective) learning rate, and therefore the stochastic noise, from decaying to zero during training. Additional components, such as batch size and data augmentation, also inject noise into the optimization process. In particular, a smaller batch size creates noisier gradient estimates since each update is computed from fewer samples, leading to more variation in the optimization path (Keskar et al., 2016; Jastrzebski et al., 2017), while data augmentation introduces extra randomness into each minibatch (Hanin & Sun, 2021). We validate these findings across architectures (MLP, Resnet, Densenet, Transformer, and GPT), tasks (SVHN, CIFAR, TinyImagenet, WILDS, and TinyStories), and settings such as transfer learning.
@@
Lastly, we study the effect of the learning rate in task arithmetic merging, which defines a different subspace of solutions than linear interpolation. We find that the geometry of the loss landscape in task arithmetic significantly changes depending on the initialization. Given a pretrained initialization (e.g. CLIP), a larger learning rate identifies solutions with greater merging compatibility (Figure 7). Moreover, the landscape is also flatter compared to a smaller learning rate. When merging solutions trained on two different downstream tasks using different learning rates, we observe that similar optimization configurations are more compatible to merge (Figure 8).
@@ 2 Preliminaries
Linear interpolation merging (Frankle et al., 2020). Linear mode connectivity refers to a phenomenon where two minima with similar performance can be connected by a linear path in the parameter space without significant performance degradation along that path. Formally, given two neural networks with parameters $\bm{\theta}_{A}$ and $\bm{\theta}_{B}$, we can define a linear interpolated model $\bm{\theta}_{li}$ as:
@@
Task arithmetic merging (Ilharco et al., 2023). Given a base model $\bm{\theta}_{base}$, a finetuned model $\bm{\theta}_{t}$ on task $t$, the task vector is defined as $\tau_{t}=\bm{\theta}_{t}-\bm{\theta}_{base}$. This vector $\tau_{t}$ encodes all the properties of the task $t$. Interestingly, task arithmetic enables operations such as addition and scaling of different task vectors, creating a merged model $\bm{\theta}_{ta}$ as:
@@
Effective noise scale. Stochastic optimization injects gradient noise $\xi$ into the optimization dynamics. Writing the minibatch gradient as $g_{t}=\nabla\mathcal{L}(\bm{\theta}_{t})+\xi_{t}$ where $\mathbb{E}[\xi_{t}]=0$ and $\mathrm{Cov}[\xi_{t}]\approx\Sigma(\bm{\theta}_{t})/B$, the stochastic update (with momentum $\mu$ and decoupled weight decay $\lambda$) can be viewed as a discretized stochastic differential equation whose “temperature” scales with learning rate and inversely with batch size (Mandt et al., 2017; Smith et al., 2018; McCandlish et al., 2018). This can be summarized by the effective noise scale:
$\mathcal{S}_{\mathrm{eff}}(\eta,B,\mu;\mathcal{A})\propto\frac{\eta}{B(1-\mu)^{2}}\,\mathrm{tr}\,\Sigma_{\mathcal{A}}(\bm{\theta}_{t}),\qquad\tilde{\mathcal{S}}=\frac{\eta}{B(1-\mu)^{2}},$ (3)
where $\eta$ is the learning rate, $B$ the batch size, and $\Sigma_{\mathcal{A}}$ the gradient-noise covariance, which increases with stronger data augmentation $\mathcal{A}$ and data diversity. We report the results using the normalized proxy $\smash{\tilde{\mathcal{S}}}$ and show how it controls the merging effectiveness.
@@ 3 The optimizer’s implicit bias on linear interpolation
This section presents our key findings for linear interpolation merging. We begin by presenting the effective noise scale as the unifying implicit bias controlling the effectiveness of merging. Then, we demonstrate that each optimizer component affects merging effectiveness through this noise.
@@ 3.1 Effective noise scale as a unifying factor
(a) Learning rate
(b) Batch size
(c) Effective noise
Figure 1: Effective noise scale controls the effectiveness of merging. The y-axis reports the test accuracy gain of merged models. On the x-axis, when plotting (a) batch sizes against learning rates or (b) vice versa, there is no clear trend. When reparameterized in terms of (c) effective noise scale, the curves are aligned, highlighting the interaction between different components for merging.
As introduced in Section 2, the effective noise scale $\smash{\tilde{\mathcal{S}}}$ captures the joint interaction of learning rate, batch size, momentum, and augmentation of the stochasticity in SGD. Rather than treating these hyperparameters as independent, we study how $\smash{\tilde{\mathcal{S}}}$ offers a unifying view on the compatibility of single models under linear interpolation merging. We use Resnet18 on CIFAR100 and sweep across different learning rates and batch sizes. The weight decay is fixed at 5e-4, and the random flip and crop augmentation are used. To evaluate the effectiveness of merging, we define performance gain as $g(\bm{\theta}_{merge},\bm{\theta}_{single})=f(\bm{\theta}_{merge})-f(\bm{\theta}_{single})$ given an evaluation function $f:\Theta\times\mathcal{D}\rightarrow\mathbb{R}$. Section A.1 describes the training and merging setup.
The results in Figure 1 illustrate how $\smash{\tilde{\mathcal{S}}}$ affects the merging effectiveness. When we vary the learning rate or batch size independently, there is no clear trend across both dimensions at the same time. For example, when increasing the learning rate for a fixed batch size to $B=16$, the accuracy gains monotonically decrease, whereas the opposite holds for a larger batch size of $B=128$ or $B=256$, which is commonly preferred in practice. However, once we represent the x-axis in terms of the normalized effective noise $\smash{\tilde{\mathcal{S}}}$, capturing both learning rate and batch size together, the different curves become aligned. Importantly, this curve is non-monotonic: merging effectiveness improves as $\smash{\tilde{\mathcal{S}}}$ increases from small values, reaches a “sweet spot”, and then degrades again once the noise grows too large.
In contrast to prior work that mainly links effective noise to properties of single solutions (Chaudhari et al., 2016; Mandt et al., 2017; Jastrzebski et al., 2017), our results show that it also governs its surrounding solutions. Specifically, $\smash{\tilde{\mathcal{S}}}$ not only biases SGD toward particular regions of the loss landscape, but also controls the compatibility of solutions found in different runs under linear interpolation merging. In the following sections, we ablate each optimizer component and analyze how it contributes to the merging effectiveness.
3.2 Large learning rate produces more compatible solutions
@@
Figure 2: Larger learning rate leads to more effective merging. (top) The test accuracy gain of all the models. (bottom) Each point represents a single model accuracy on the $x$-axis and its accuracy gain after merging on the $y$-axis. The opacity indicates the number of training epochs. For each setup, we observe that a larger learning rates have a higher accuracy gain, even when there is a smaller learning rate with equivalent single model accuracy. Note, however, solutions found using a “too large” learning rate fail to merge (details in Section B.6).
We present empirical results for deep neural networks on vision tasks trained from scratch. The architectures used are a simple MLP, Resnet18, Densenet121 trained on SVHN, CIFAR10, CIFAR100, and TinyImageNet. Weight decay is fixed at 5e-4, batch size of 128, and random flip and crop augmentation are used across the setups. Section A.1 describes the training and merging setup.
The results in Figure 2 show the key finding of our work. Using linear interpolation with a fixed $\alpha=0.5$, we merge two solutions with the same learning rates at each checkpoint. We observe that the solutions identified with a larger learning rate are consistently more compatible to merge than those of a smaller learning rate. For example, in CIFAR100, the solutions found using an $\text{lr}=0.2$ achieve $1.2\%$ of the median gain compared to a $0.2\%$ gain of $\text{lr}=0.01$, despite having a similar performance for the single model of $\approx 75\%$ (x-axis). The same phenomenon is observed across different datasets (SVHN, CIFAR, and TinyImagenet) and architectures (MLP, Resnet, and Densenet). Furthermore, we also ensure that all the solutions are well-converged by asserting that the training loss is near-zero (details in Section B.5). As argued by Pascanu et al. (2025), it is important to understand how the implicit bias of the optimizer alters the final solutions and how to leverage this bias to find better solutions. Furthermore, recent works found different implicit biases of training with a larger learning rate, such as a sparser activation (Andriushchenko et al., 2023b; Sadrtdinov et al., 2024), a different order for feature learning order (Li et al., 2019), and a flatter solution (Andriushchenko et al., 2023a). Our results demonstrate an additional benefit: larger learning rate has an implicit bias on the loss landscape, identifying more compatible solutions for model merging. Note, however, Appendix Figure 15 shows that a too large learning rate leads to instabilities or failures in model merging.
@@ 3.3 Weight decay and effective learning rate
Figure 3: Weight decay has a similar effect as the learning rate. For CIFAR100 and TinyImagenet, we use scale-invariant networks (w/ normalization layers) and observe that a larger weight decay can not only improve the accuracy of the single model, but also the accuracy gain via the effective learning rate (Van Laarhoven, 2017). For MLP trained on SVHN, there is no trend as the architecture is not scale-invariant.
[…]
@@
We hypothesize that weight decay has a similar effect to the learning rate in model merging. That is, solutions found with a larger weight decay are easier to merge than those found with a smaller or no weight decay. Note that this should hold only for scale-invariant networks. We use the same experimental setup as in Section 3.2, except that we now sweep across different weight decay values instead of learning rates.
@@
The results in Figure 3 confirm our hypothesis about the implicit bias of weight decay. Larger weight decay increases the effective learning rate for scale-invariant networks during training, affecting also the model merging effectiveness. For example, in TinyImagenet, the solutions found using a $\text{wd}=0.0005$ report $1.2\%$ of the median gain compared to a $0.5\%$ of other wd values. The same trend is observed in other settings. Furthermore, interestingly, for the architecture MLP trained on SVHN, larger weight decay does not differ from smaller ones in terms of merging effectiveness. This validates that weight decay affects model merging only in scale-invariant architectures. Our results extend how the weight decay affects the loss landscape of an individual minima (Van Laarhoven, 2017; D’Angelo et al., 2024) to its connection with other minima. Lastly, similarly to the learning rate, Appendix Figure 16 shows that a too large weight decay leads to failure in model performance and model merging.
3.4 Batch size and data augmentation
(d) Batch size
(e) Data augmentation

### v2

@@ Abstract
Model merging combines independent solutions with different capabilities into a single one while maintaining the same inference cost. Two popular approaches are linear interpolation, which simply averages multiple model weights, and task arithmetic, which combines task vectors obtained by the difference between finetuned and base models. While useful in practice, what properties make merging effective are poorly understood. This paper explores how the optimization dynamics affect the loss landscape geometry and its impact on merging success. We show that a single quantity – the effective noise scale – unifies the impact of different optimizer components on model merging. Across architectures and datasets, merging success is a non-monotonic function of the effective noise scale, with a distinct optimum. Decomposing this quantity, we find that larger learning rates, stronger weight decay, smaller batch sizes, and data augmentation all independently modulate the effective noise scale and exhibit the same qualitative trend. Unlike prior work connecting optimizer noise to the flatness or generalization of individual minima, we show that it also affects the global loss landscape, predicting when independently trained solutions can be successfully merged. Our findings broaden the understanding of how optimization shapes the loss landscape geometry and its consequences for model merging, suggesting that training dynamics could be further manipulated to improve model merging.
@@ 1 Introduction
Model merging combines independently trained solutions to fuse their unique capabilities into a single solution. These methods have been applied to either further improve the performance of a single model (Wortsman et al., 2022) or to combine models with different, but similar, capabilities (Ilharco et al., 2023). Notably, this is achieved without additional computational cost at inference time. Given these practical advantages, merging methods have been applied to state-of-the-art architectures (Yadav et al., 2024; Cohere et al., 2025). However, these approaches rely on extensive trial-and-error, requiring practitioners to train and evaluate multiple models to identify which candidates successfully merge. A fundamental unresolved question is why some models are compatible for merging while others fail, despite achieving similar individual performance.
@@
The success of model merging depends on the mode connectivity phenomenon, the geometric property that independent solutions can be connected by paths of low-loss models (Draxler et al., 2018; Garipov et al., 2018)Frankle et al. (2020) demonstrated a stricter condition, linear mode connectivity, showing that solutions sharing a common initial optimization trajectory can be even connected by a linear path with similar loss. This finding suggests that optimization dynamics, not just final convergence, fundamentally shape the geometry of the loss landscape between solutions. Yet, despite this insight, well-studied factors that affect the optimization dynamics, such as learning rate, weight decay, or batch size, are not well understood for the model merging loss landscape.
@@
In this work, we study the role of optimization dynamics on the outcome of model merging. First, we show that different optimizer components (learning rate, weight decay, batch size, and data augmentation) control the same underlying factor, the effective noise scale, which captures this magnitude of stochastic fluctuations in the optimization trajectory. Our experiments demonstrate how the noise magnitude determines whether independently trained solutions occupy merging-compatible regions of the loss landscape. A noise regime that is too small or too large leads to almost no merging benefits, whereas a moderately large noise regime maximizes the benefits. We decompose the noise into individual components, showing that each one exhibits the same qualitative trend. Starting from the learning rate, we find that a larger and stable value can consistently identify solutions that merge more effectively than those trained with smaller learning rates, even when both converge to similar generalization performance on the test data. Since decaying any learning rate during the optimization can always lead to a converged solution, it is perhaps surprising that simply starting with a larger learning rate can change the merging outcomes. However, beyond classical research showing direct advantage of large learning rates on generalization (Keskar et al., 2016), recent works presented different implicit biases of larger learning rate, such as a sparser activation (Andriushchenko et al., 2023b), a different sequence of pattern learning (Li et al., 2019), and a flatter solution (Andriushchenko et al., 2023a). Our results extend these benefits beyond single-task performance, showing that a larger learning rate unlocks more effective merging.
@@
Similarly, we find that a large weight decay also enables more effective merging, beyond improvements in the single model performance. This can be explained with the notion of the effective learning rate (Van Laarhoven, 2017; Hoffer et al., 2018), which suggests that the main role of weight decay is to prevent the (effective) learning rate, and therefore the stochastic noise, from decaying to zero during training. Additional components, such as batch size, momentum, and data augmentation, also inject noise into the optimization dynamics. A smaller batch size creates noisier gradient estimates since each update is computed from fewer samples, leading to more variation in the optimization path (Keskar et al., 2016; Jastrzebski et al., 2017). Higher momentum smooths gradient updates and reduces oscillations, biasing the optimizer toward broader, flatter regions of the loss landscape (Sutskever et al., 2013). Data augmentation introduces extra randomness into each minibatch (Hanin and Sun, 2021). We validate these findings across different architectures, tasks, permutation symmetries, and in transfer learning setups.
@@
Lastly, we extend the analysis to task arithmetic merging, which defines a different subspace of solutions than simple linear interpolation. We find that the geometry of the loss landscape in task arithmetic significantly changes depending on the initialization. Given a pretrained initialization (e.g. CLIP), a larger learning rate identifies solutions with easier merging compatibility (Figure 7). Moreover, the landscape is also flatter compared to a smaller learning rate. Finally, we find that merging solutions trained on different tasks with a moderately larger noise has the best performance but at the cost of losing compatibility with models trained with different noise (Figure 8).
@@ 2 Preliminaries
Linear interpolation merging (Frankle et al., 2020). Linear mode connectivity refers to a phenomenon where two minima with similar performance can be connected by a linear path in the parameter space without significant performance degradation along that path. Formally, given two models with parameters $\bm{\theta}_{A}$ and $\bm{\theta}_{B}$, we can define a linear interpolated model $\bm{\theta}_{li}$ as:
@@
Task arithmetic merging (Ilharco et al., 2023). Given a base model $\bm{\theta}_{base}$, a finetuned model $\bm{\theta}_{t}$ on task $t$, the task vector is defined as $\tau_{t}=\bm{\theta}_{t}-\bm{\theta}_{base}$. The task vector $\tau_{t}$ captures the parameter change induced by fine-tuning on task $t$. Interestingly, task arithmetic enables operations such as addition and scaling of different task vectors, creating a merged model $\bm{\theta}_{ta}$ as:
@@
Effective noise scale. Stochastic optimization introduces randomness through minibatch sampling. Writing the minibatch gradient as $g_{t}=\nabla\mathcal{L}(\bm{\theta}_{t})+\xi_{t}$ with $\mathbb{E}[\xi_{t}]=0$ and $\mathrm{Cov}[\xi_{t}]\approx\Sigma_{\mathcal{A}}(\bm{\theta}_{t})/B$, the SGD update can be viewed as a discretized SDE whose diffusion strength scales proportionally with the learning rate $\eta$ and inversely with the batch size $B$ (Mandt et al., 2017). In the standard momentum parameterization, this diffusion is further rescaled by $(1-\mu)^{-1}$ (Smith et al., 2018). The remaining magnitude is task and data-dependent through the gradient-noise covariance $\mathrm{tr}\,\Sigma$ (McCandlish et al., 2018). We summarize all these effects by the effective noise scale:
$\mathcal{S}_{\mathrm{eff}}(\eta,B,\mu;\mathcal{A})\propto\frac{\eta}{B(1-\mu)}\,\mathrm{tr}\,\Sigma_{\mathcal{A}}(\bm{\theta}_{t}),\qquad\tilde{\mathcal{S}}=\frac{\eta}{B(1-\mu)}.$ (3)
Note that $\Sigma_{\mathcal{A}}$ depends on data augmentation and diversity $\mathcal{A}$. When $\mathcal{A}$ is fixed across runs, we can treat $\mathrm{tr}\,\Sigma_{\mathcal{A}}(\bm{\theta}_{t})$ as approximately constant, so $\tilde{\mathcal{S}}$ provides a practical proxy for comparing noise levels across runs.
@@ 3 The optimizer’s implicit bias on linear interpolation
This section presents our key findings for linear interpolation merging. We begin by presenting the effective noise scale as the unifying implicit bias controlling model merging. Then, we demonstrate that each optimizer component affects merging effectiveness through this noise.
@@ 3.1 Effective noise scale as a unifying factor
Figure 1: Effective noise scale controls the effectiveness of merging. The $y$-axis reports the test accuracy gain of merged models over single models. In (left) and (middle), plotting batch size against learning rate (or vice versa) reveals no clear trend. When reparameterized in terms of (right) effective noise scale, the joint interaction between different components emerges.
As introduced in Section 2, the effective noise scale $\smash{\tilde{\mathcal{S}}}$ captures the joint interaction of learning rate, batch size, momentum, and augmentation of the optimization dynamics. Rather than treating these hyperparameters as independent, we present how $\smash{\tilde{\mathcal{S}}}$ offers a unifying view on the compatibility of single models under linear interpolation merging. We train ResNet18 on CIFAR100 using SGD, sweeping jointly the learning rate $\eta$ and batch size $B$. The weight decay $\lambda=5e-4$ is fixed, and random flip and crop augmentations are used. To evaluate the merging effectiveness, we define performance gain as $g(\bm{\theta}_{merge},\bm{\theta}_{single})=f(\bm{\theta}_{merge})-f(\bm{\theta}_{single})$ given an evaluation function $f:\Theta\times\mathcal{D}\rightarrow\mathbb{R}$. Section A.1 describes the training and merging setup.
Figure 1 illustrates how $\smash{\tilde{\mathcal{S}}}$ affects the merging effectiveness. When the learning rate or batch size are varied independently, there is no clear trend across both dimensions. For example, when increasing the learning rate for a fixed batch size to $B=16$, the accuracy gains monotonically decrease, whereas the opposite holds for a larger batch size of $B=128$ or $B=256$, which is commonly preferred in practice. However, once we reparametrize the $x$-axis in terms of $\smash{\tilde{\mathcal{S}}}$, capturing both learning rate $\eta$ and batch size $B$ together, the different curves become aligned. Importantly, this curve is non-monotonic: merging effectiveness improves as $\smash{\tilde{\mathcal{S}}}$ increases from small values, reaches a “critical point”, and then degrades again once the noise grows too large. Appendix Figure 14 presents the same trend for DenseNet121 trained on TinyImageNet.
In contrast to prior work that mainly links effective noise to properties of single solutions (Chaudhari et al., 2016; Mandt et al., 2017; Jastrzebski et al., 2017), our results show that it also governs its surrounding solutions. Specifically, $\smash{\tilde{\mathcal{S}}}$ not only biases the optimization dynamics toward particular regions of the loss landscape, but also controls the compatibility of solutions found in different runs under linear interpolation merging. In the following sections, we control each optimizer component in isolation to analyze how it contributes to the merging effectiveness.
3.2 Large learning rate identifies more compatible solutions
@@
Figure 2: Larger learning rate leads to more effective merging. (top) The test accuracy gain of all the models. (bottom) Each point represents a single model performance $\theta_{A}$ on the $x$-axis and its accuracy gain after merging on the $y$-axis. The opacity indicates the number of training epochs. For each setup, we observe that larger learning rates have a higher accuracy gain, even when there is a smaller learning rate with equivalent accuracy on the $x$-axis. Note, however, solutions with a “too large” learning rate fail to merge (Section B.7).
We present empirical results for vision tasks trained from scratch using a simple MLP, ResNet-18, and DenseNet121 on SVHN, CIFAR-10, CIFAR-100, and TinyImageNet with SGD. We fix $\lambda=5e-4$, $B=128$, $\mu=0.9$, and random flip and crop augmentation are used across the setups. We ensure that all the solutions are well-converged by verifying that the training loss is near-zero. Section A.1 describes the training and merging setup.
Under linear interpolation with a fixed $\alpha=0.5$, we merge two solutions trained with the same learning rates $\eta$. Figure 2 shows that the solutions identified with a larger $\eta$ are consistently more compatible to merge than those found with smaller values. For example, in CIFAR100, the solutions found using an $\eta=2e-1$ achieve $1.2\%$ of the median gain compared to a $0.2\%$ gain of $\eta=1e-2$, despite having a similar single model performance of $\approx 75\%$ ($x$-axis). This demonstrates that larger noise $\eta$ induces a more effective loss landscape geometry for model merging. The same phenomenon is observed across different datasets (SVHN, CIFAR, and TinyImageNet) and architectures (MLP, ResNet, and DenseNet).
Pascanu et al. (2025) argue that understanding how the implicit bias of an optimizer shapes the final solution, and how to leverage this bias to find better models, is an important problem. We address this directly by showing how $\eta$ biases the model merging process. Additionally, recent works found different implicit biases of training with a large $\eta$, such as a sparser activation (Andriushchenko et al., 2023b; Sadrtdinov et al., 2024), a different feature-learning order (Li et al., 2019), or a flatter solution (Andriushchenko et al., 2023a). Our results demonstrate an additional benefit: larger learning rate has an implicit bias on the loss landscape, identifying more compatible solutions for model merging up to a certain point (see Appendix Figure 16).
@@ 3.3 Weight decay and effective learning rate
Figure 3: Weight decay $\lambda$ has a similar effect as the learning rate $\eta$ for model merging. For CIFAR100 and TinyImageNet, we use scale-invariant networks (w/ normalization layers) and observe that a larger weight decay can not only improve the accuracy of the single model, but also the accuracy gain via the effective learning rate (Van Laarhoven, 2017). For the MLP architecture, there is no trend as it is not scale-invariant.
The traditional understanding of the role of weight decay regularization is that it reduces overfitting by proportionally decaying the weights towards zero, favouring simpler models. In practice, this is achieved by adding a penalty term $\lambda\|\bm{\theta}\|^{2}_{2}$ to the objective $\mathcal{L}(\bm{\theta})$. However, modern neural network architectures ubiquitously use normalization layers (Ioffe and Szegedy, 2015; Ba et al., 2016) and are therefore weight scale-invariant as $f(\bm{x},\alpha\bm{\theta})=f(\bm{x},\bm{\theta})$. Then, what is the new role of weight decay regularization in scale-invariant networks? Van Laarhoven (2017); Hoffer et al. (2018) prove and demonstrate that weight decay controls the effective learning rate, and thus the optimization dynamics during training. In practice, for a scale-invariant network with $\lambda=0$, its gradient norm decays proportionally to the increase of weight norm $\|\nabla\mathcal{L}\|^{2}_{2}\propto 1/\|\bm{\theta}\|^{2}_{2}$, which leads to the vanishing of the effective learning rate (Van Laarhoven, 2017).
@@
We hypothesize that weight decay $\lambda$ has a similar effect to the learning rate $\eta$ for model merging. That is, solutions identified with a larger weight decay are easier to merge than those found using a smaller or no weight decay. Note that this holds only for scale-invariant networks, since the $\|\bm{\theta}\|^{2}_{2}$ of a non scale-invariant network is already constrained by the loss without the weight decay penalty (Van Laarhoven, 2017). We use the same experimental setup as in Section 3.2, except that we now sweep across different weight decay values instead of learning rates.
@@
Figure 3 confirms our hypothesis about the implicit bias of weight decay. Larger $\lambda$ increases the effective learning rate for scale-invariant networks during training, affecting the model merging effectiveness. For example, in TinyImageNet, the solutions found using a $\lambda=5e-4$ report $1.2\%$ of the median gain compared to a $0.5\%$ of other $\lambda$ values. The same trend is observed in multiple settings. Furthermore, for non scale-invariant architecture MLP trained on SVHN, different values of $\lambda$ regularization have similar impact on the merging effectiveness, confirming that weight decay affects primarily when merging scale-invariant architectures. Our results extend how weight decay affects the loss landscape of an individual minimum (Van Laarhoven, 2017; D’Angelo et al., 2024) to its connection with other minima. Similar to the learning rate, Appendix Figure 17 shows that excessive weight decay leads to failure in model performance and model merging.
3.4 Batch size, momentum, and data augmentation
(a) Batch size ($B$)
(b) Momentum ($\mu$)
(c) Augmentation ($\Sigma_{\mathcal{A}}$)